# Weighted Covariance Intersection for Range-based Distributed Cooperative Localization of Multi-Agent Systems

### v1

@@ III-A State Prediction Phase
The state prediction phase includes INS mechanism and error propagation, which have been extensively studied and are not the primary focus of our research. Therefore, we will not delve into these aspects here; readers are referred to [25, 23] for further details. It is worth mentioning that during the state prediction phase, states other than position, such as velocity, attitude, and IMU biases, affect the position estimation results. Therefore, for 3D cooperative localization, we must not only focus on the accuracy of position estimation but also ensure a balanced and satisfactory estimation of all states.
Figure 2: Range-based distributed cooperative localization framework for 3D scenarios. Here $\{\boldsymbol{d}_{i}^{\mathcal{U}}\}$ and $\{\boldsymbol{d}_{i}^{\mathcal{A}}\}$ denote the sets of distance measurements from agent $i$ to other agents and anchors, respectively. $\text{info}_{i}$ and $\{\text{info}_{j}\}$ represent the information of agent $i$ that needs to transmitted to other agents, and the information received from other agents, respectively.
@@ III-B1 Observation Equation
As for the state update phase, we begin with the observation equations for distance measurements. Suppose the predicted state of agent $i$ is defined as follows:
@@
Then, the predicted distance between agent $i$ and agent $j$, as well as the predicted distance between agent $i$ and anchor $m$ can be expressed as:
${\hat{d}_{ij}^{\mathcal{U}}}=$ $\left\lVert\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{R}{}{\hat{\boldsymbol{p}}_{j}^{n}}\right\rVert$ (7)
$\approx$ $\left\lVert\prescript{R}{}{\boldsymbol{p}_{i}^{n}}-\prescript{R}{}{\boldsymbol{p}_{j}^{n}}\right\rVert+(\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}})^{\mathsf{T}}\left(\prescript{I}{}{\delta{\boldsymbol{p}}_{i}^{n}}-\prescript{I}{}{\delta{\boldsymbol{p}}_{j}^{n}}\right)$
$+$ $(\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}})^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right]\boldsymbol{\phi}_{i}^{n}-(\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}})^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{j}}^{n}\boldsymbol{l}^{b}\right)\times\right]\boldsymbol{\phi}_{j}^{n}$
@@
where $\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}}$ and $\hat{\boldsymbol{r}}_{im}^{\mathcal{A}}$ denote the estimated unit direction vectors pointing from the positions of agent $j$’s and anchor $m$’s ranging modules to the position of agent $i$’s ranging module, respectively. Their explicit expressions are as follows:
$\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}}=\frac{\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{R}{}{\hat{\boldsymbol{p}}_{j}^{n}}}{\left\lVert\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{R}{}{\hat{\boldsymbol{p}}_{j}^{n}}\right\rVert}\qquad\hat{\boldsymbol{r}}_{im}^{\mathcal{A}}=\frac{\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{A}{}{{\boldsymbol{p}}_{m}^{n}}}{\left\lVert\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{A}{}{{\boldsymbol{p}}_{m}^{n}}\right\rVert}$ (8)
Since we use the error-state formulation for error propagation and state updates, we apply differential measurements rather than the original measurements during the state update phase. The differential measurements, i.e., the differences between the predicted distances and actual distance measurements, can be expressed as:
$\delta{d_{ij}^{\mathcal{U}}}$ $={\hat{d}_{ij}^{\mathcal{U}}}-{\tilde{d}_{ij}^{\mathcal{U}}}=\boldsymbol{H}_{ij}^{\mathcal{U}}\delta\boldsymbol{x}_{i}+\boldsymbol{G}_{ij}^{\mathcal{U}}\boldsymbol{v}_{ij}^{\mathcal{U}}$ (9)
@@
Here, $\boldsymbol{v}_{ij}^{\mathcal{U}}$ and $v_{im}^{\mathcal{A}}$ are the noise of the differential measurements, where $v_{im}^{\mathcal{A}}=-n_{d_{im}^{\mathcal{A}}}$, and the expressions of $\boldsymbol{v}_{ij}^{\mathcal{U}}$ and its corresponding covariance matrix $\boldsymbol{R}_{v_{ij}}$ are given by
$\boldsymbol{v}_{ij}^{\mathcal{U}}=\left[\left(\prescript{I}{}{\delta{\boldsymbol{p}}_{j}^{n}}\right)^{\mathsf{T}},\left(\boldsymbol{\phi}_{j}^{n}\right)^{\mathsf{T}},n_{d_{ij}^{\mathcal{U}}}\right]^{\mathsf{T}}$ (10)
$\boldsymbol{R}_{ij}^{\mathcal{U}}=\begin{bmatrix}&\boldsymbol{P}_{{p}_{j}},&\boldsymbol{P}_{{p}_{j}{\phi}_{j}},&\boldsymbol{0}_{3\times 1}\\ &\boldsymbol{P}_{{p}_{j}{\phi}_{j}}^{\mathsf{T}},&\boldsymbol{P}_{{\phi}_{j}},&\boldsymbol{0}_{3\times 1}\\ &\boldsymbol{0}_{1\times 3},&\boldsymbol{0}_{1\times 3},&\sigma_{r}^{2}\end{bmatrix}$ (11)
Here $\boldsymbol{P}_{{p}_{j}}$, $\boldsymbol{P}_{{\phi}_{j}}$, $\boldsymbol{P}_{{p}_{j}{\phi}_{j}}$ are the submatrices of the state covariance matrix of agent $j$, corresponding to the covariance matrix of its position, attitude, and their cross-covariance, respectively. Additionally, the expressions of $\boldsymbol{H}_{ij}^{\mathcal{U}}$, $\boldsymbol{G}_{ij}^{\mathcal{U}}$ and $\boldsymbol{H}_{ij}^{\mathcal{A}}$ are given by
$\boldsymbol{H}_{ij}^{\mathcal{U}}$ $=\left[\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}}\right)^{\mathsf{T}},\boldsymbol{0}_{3\times 3},\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}}\right)^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right],\boldsymbol{0}_{3\times 6}\right]$ (12)
$\boldsymbol{G}_{ij}^{\mathcal{U}}$ $=\left[-\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}}\right)^{\mathsf{T}},-\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}}\right)^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right],-1\right]$
@@
Assume that agent $i$ obtains the distances to $M$ anchors and $N$ other agents for state updates at time step $k$. Without loss of generality, let the indices of the $M$ anchors be $1$ to $M$, and the indices of the $N$ other agents be $1$ to $N$. Let the measurement vector formed by the differential distances to the $M$ anchors be denoted as $\boldsymbol{z}_{i}^{\mathcal{A}}=\left[\delta d_{i1}^{\mathcal{A}},\cdots,\delta d_{iM}^{\mathcal{A}}\right]^{\mathsf{T}}\in\mathbb{R}^{M\times 1}$, with the corresponding covariance matrix being $\boldsymbol{R}_{i}^{\mathcal{A}}=\sigma_{r}^{2}\boldsymbol{I}_{M}$; Let the differential distances to the $N$ other agents and their corresponding covariance matrices be denoted as $\left\{\delta d_{ij}^{\mathcal{U}},\boldsymbol{R}_{ij}^{\mathcal{U}}\right\}$, $j=1,\ldots,N$, respectively. The predicted state and the error-state covariance matrix of agent $i$ before the update are given by $\left\{\hat{\boldsymbol{x}}_{i,k|k-1},\boldsymbol{P}_{i,k|k-1}\right\}$.
@@ III-B2 Independent Update
The distance measurements between agent $i$ and anchors, which are independent of the agent’s state, can be directly used for state update via EKF. Since the distances to the $M$ anchors are mutually independent, there is no difference between updating them sequentially as individual observations or combining them into a single observation vector for concurrent updating. We choose the latter approach, where all anchor measurements are aggregated into a single observation vector for the update. The corresponding update procedure can be written as:
@@ III-B3 Correlated Update
For distance measurements between agents, consider the measurement between agent $i$ and agent $j$ as an example, as shown in (9) and (10). The differential distance $\delta d_{ij}^{\mathcal{U}}$ is used for updating agent $i$’s state, and its noise includes the state estimation error of agent $j$. This leads to potential correlations between the differential distance measurement and agent $i$’s state. In a distributed framework, tracking the state correlation between agent $i$ and agent $j$ is extremely challenging or even infeasible. If the correlation is ignored and assumed to be independent, it could lead to the “data incest” problem [26], resulting in state estimation divergence. Therefore, our work focuses on this correlated update process, with the goal of achieving effective state updates using relative distance measurements between agents under unknown correlations, ensuring accurate and consistent state estimation in 3D cooperative localization scenarios.
IV Methodology
In this section, we present our methods for DCL in 3D scenarios. We begin by revisiting CI and discussing three primary drawbacks of the classical CI criteria in recursive filtering. Next, we introduce a weighting mechanism, namely WCI, to mitigate the negative impacts of CI and analyze its properties. Finally, we propose a WCI-based approach for correlated updates in DCL, which accounts for the characteristics of CI fusion and the error propagation rule of INS, balancing the estimation of all state components with scale disparity.
@@ IV-A1 Formulas of Covariance Intersection
Typically, the cost function is chosen as the trace or determinant of the fused covariance matrix, i.e., $f(\boldsymbol{P})=\operatorname{tr}(\boldsymbol{P})$ or $\text{det}(\boldsymbol{P})$; we refer to them as the classical criteria.
@@
IV-A2 Drawbacks of the Classical Criteria
Although the classical CI criteria have achieved widespread success in various applications, they can degrade performance in certain scenarios. We identify three primary drawbacks of classical criteria in the context of recursive filtering:
  • Scale imbalance
    When state components differ substantially in units or magnitudes, the trace-based criterion becomes dominated by the larger-scale components and tends to overly magnify the uncertainty of small-scale components. We give a simple two-dimensional example to demonstrate it.
    Let $\boldsymbol{P}_{1}=\begin{bmatrix}1&0\\ 0&0.01\end{bmatrix}$ and $\boldsymbol{P}_{2}=\begin{bmatrix}1.09&0\\ 0&0.001\end{bmatrix}$. Although $\boldsymbol{P}_{2}$ reduces the second component’s variance by 90% dramatically, its 9% increase in the first component drives the trace-minimization fusion (CI-trace) to favor $\boldsymbol{P}_{1}$. As shown in Fig. 3(a), the fused covariance matrix thus remains almost identical to $\boldsymbol{P}_{1}$, offering negligible improvement for the small-scale component. By contrast, determinant-based criterion (CI-det) favors $\boldsymbol{P}_{2}$ because the determinant corresponds to the uncertainty volume in which components with small scales can play a role. However, both criteria lack the ability to flexibly adjust and balance the estimation performance of different components in the presence of significant scale disparity.
  • Correlation mismatch
    The actual covariance matrix of the CI estimate, given by (18), depends on the actual cross‐covariance $\boldsymbol{P}_{12}^{*}$, whereas the nominal covariance matrix $\boldsymbol{P}$ in (15) is completely independent of cross‐correlations. This ensures consistency by guaranteeing that the nominal covariance matrix dominates the actual covariance matrix in Löwner partial order for all admissible cross-correlations, but offers no alignment of intra-state correlations. When the actual correlations diverge substantially from those implied by $\boldsymbol{P}$, recursive filtering can experience severe error divergence, particularly under partial observation or scale disparity cases.
    $\begin{split}\boldsymbol{P}^{*}&=\mathbb{E}\left\{\delta{\boldsymbol{x}}\delta{\boldsymbol{x}}^{\mathsf{T}}\right\}\\ &=\boldsymbol{P}[\ \omega^{2}\boldsymbol{P}_{1}^{-1}\boldsymbol{P}_{1}^{*}\boldsymbol{P}_{1}^{-1}+(1-\omega)\omega\boldsymbol{P}_{1}^{-1}\boldsymbol{P}_{12}^{*}\boldsymbol{P}_{2}^{-1}+\\ &(1-\omega)\omega\boldsymbol{P}_{2}^{-1}\boldsymbol{P}_{12}^{*\mathsf{T}}\boldsymbol{P}_{1}^{-1}+(1-\omega)^{2}\boldsymbol{P}_{2}^{-1}\boldsymbol{P}_{2}^{*}\boldsymbol{P}_{2}^{-1}\ ]\boldsymbol{P}\end{split}$ (18)
    For example, consider two estimates of $\boldsymbol{x}$: the first estimate is full observed with $\boldsymbol{P}_{1}=\begin{bmatrix}1&0.008\\ 0.008&0.0001\end{bmatrix}$, while the second estimate is partially observed with $\boldsymbol{H}=[1,0]$ and $\boldsymbol{P}_{2}=0.25$. Without loss of generality, we assume that the two estimates are uncorrelated—though this is unknown a priori—so we apply CI to fuse them. Fig. 3(b) shows that the CI‐fused nominal covariance ellipse encloses the actual covariance ellipse, confirming CI’s consistency and conservativeness. However, the nominal correlation coefficient given by CI-trace, $\rho_{1}=0.08$, deviates drastically from the actual correlation coefficient, $\rho_{1}^{*}=0.55$, which can lead to unsatisfactory updates in subsequent recursive filtering. In contrast, the correlation deviation of CI-det is relatively small but still exists in this example, suggesting that it is feasible to select an appropriate optimization criterion to mitigate the correlation mismatch problem.
(a) Example for scale imbalance
(b) Example for correlation mismatch
Figure 3: Covariance ellipses before and after CI fusion (axes scaled differently for better demonstration). (a) The covariance ellipse of CI-trace almost overlaps that of $\boldsymbol{P}_{1}$, showing that fusion is dominated by the larger-scale state component—so the fused estimate closely matches the lower-uncertainty estimate along that dimension. While CI-det can account for smaller-scale state component, but still lacks balance between two components. (b) The nominal covariance ellipse encloses the actual covariance ellipse but differs markedly in shape and orientation, reflecting a significant mismatch between implied and actual inter-component correlations.
  • Lack of task relevance
    As Hurley pointed out, the classical determinant-based criterion corresponds to minimizing the Shannon information of the fused density function [28]. However, the perspective of information theory offers little interpretation of performance in physical world. Meanwhile, the classical trace-based criterion implements a minimum MSE objective. However, when the state vector comprises heterogeneous components (e.g., position versus orientation), the aggregated MSE lacks a clear physical interpretation as well. In summary, both the two classical criteria ignores the task‐dependent significance of individual components, rendering them too coarse for applications that require prioritization among state variables.
IV-B Weighted Covariance Intersection
The classical trace-based criterion is derived from the principle of minimizing MSE of the estimate—equivalently, the trace of the actual covariance matrix (cf. (19)). In practice, since the actual covariance matrix is unavailable, the trace of the nominal covariance matrix is minimized instead.
@@ IV-B Weighted Covariance Intersection
As noted above, minimizing MSE alone is too coarse and lacks task specificity. To address this, Arambel et al. introduced relative weights to control the estimation process [15], termed WCI. However, they treated these weights merely as tuning knobs to improve the accuracy of a chosen component, without providing a theoretical rationale. We show in (20) that minimizing the weighted trace of the fused covariance matrix is mathematically equivalent to minimizing a WMSE over the entire state. By selecting the weighting matrix $\boldsymbol{W}$ to reflect task‐dependent priorities or target WMSE levels, WCI reintroduces task relevance to the fusion criterion.
@@
Moreover, by assigning greater weight to smaller-scale components, the weighting mechanism helps mitigate scale imbalance and ensures a more balanced fusion across all state variables.
Last but not least, WCI provides additional flexibility in selecting the fusion weight $\omega$, which may help mitigate correlation mismatch. Although the actual intra‐state correlations are unpredictable due to the unknown cross-correlation in (18), in the limiting cases $\omega=0$ or $1$ the cross‐correlation term vanishes from the fusion equations, ensuring that the unknown cross-correlations do not affect the final fused estimate. Thus, selecting $\omega$ closer to 0 or $1$ may reduce the potential deviation between the actual covariance matrix and its nominal counterpart.
In addition, we identify two properties of WCI: consistency and optimality, which are summarized in Proposition 1 and Proposition 2, respectively.
Proposition 1.
(Consistency) For any weight matrix, WCI provides consistent estimates.
Proof.
Because CI provides consistent estimates for any $\omega\in[0,1]$, and WCI only changes the selection of $\omega$ which is still within $[0,1]$, the resulting estimates remain consistent. ∎
Proposition 2.
(Optimality) For any positive-definite weight matrix, WCI provides the optimal bound for the fusion of two estimates under unknown correlations in terms of WMSE.
The proof of Proposition 2 is provided in Appendix.
In summary, WCI effectively mitigates the three primary drawbacks of the classical criteria of CI in recursive filtering. By selecting task‐specific weighting matrices, practitioners can tailor CI to their needs, potentially achieving improved estimation performance.
IV-C WCI-based Correlated Update in DCL

### v2

@@ III-A State Prediction Phase
The state prediction phase includes INS mechanism and error propagation, which has been extensively studied and are not the primary focus of our research. Therefore, we will not discuss these aspects here; readers are referred to [29, 27] for further details. It is worth mentioning that during the state prediction phase, states other than position, such as velocity, attitude, and IMU biases, affect the position estimation results. Therefore, for 3D cooperative localization, we must not only focus on the accuracy of position estimation but also ensure a balanced and satisfactory estimation of all state components.
Figure 3: Range-based distributed cooperative localization framework for 3D scenarios. Here $\{\boldsymbol{d}_{i}^{\mathcal{V}}\}$ and $\{\boldsymbol{d}_{i}^{\mathcal{A}}\}$ denote the sets of distance measurements from vehicle $i$ to other vehicles and anchors, respectively. $\text{info}_{i}$ and $\{\text{info}_{j}\}$ represent the information of vehicle $i$ that needs to transmitted to other vehicles, and the information received from other vehicles, respectively.
@@ III-B1 Observation Equation
As for the state update phase, we begin with the observation equations for distance measurements. Suppose the predicted state of vehicle $i$ is defined as follows:
@@
Then, the predicted distance between vehicle $i$ and vehicle $j$, as well as the predicted distance between vehicle $i$ and anchor $m$ can be expressed as:
${\hat{d}_{ij}^{\mathcal{V}}}=$ $\left\lVert\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{R}{}{\hat{\boldsymbol{p}}_{j}^{n}}\right\rVert$ (7)
$\approx$ $\left\lVert\prescript{R}{}{\boldsymbol{p}_{i}^{n}}-\prescript{R}{}{\boldsymbol{p}_{j}^{n}}\right\rVert+(\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}})^{\mathsf{T}}\left(\prescript{I}{}{\delta{\boldsymbol{p}}_{i}^{n}}-\prescript{I}{}{\delta{\boldsymbol{p}}_{j}^{n}}\right)$
$+$ $(\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}})^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right]\boldsymbol{\phi}_{i}^{n}-(\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}})^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{j}}^{n}\boldsymbol{l}^{b}\right)\times\right]\boldsymbol{\phi}_{j}^{n}$
@@
where $\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}}$ and $\hat{\boldsymbol{r}}_{im}^{\mathcal{A}}$ denote the estimated unit direction vectors pointing from the positions of vehicle $j$’s and anchor $m$’s ranging modules to the position of vehicle $i$’s ranging module, respectively. Their explicit expressions are as follows:
$\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}}=\frac{\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{R}{}{\hat{\boldsymbol{p}}_{j}^{n}}}{\left\lVert\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{R}{}{\hat{\boldsymbol{p}}_{j}^{n}}\right\rVert}\qquad\hat{\boldsymbol{r}}_{im}^{\mathcal{A}}=\frac{\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{A}{}{{\boldsymbol{p}}_{m}^{n}}}{\left\lVert\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{A}{}{{\boldsymbol{p}}_{m}^{n}}\right\rVert}$ (8)
Since we use the error-state formulation for error propagation and state updates, we apply differential measurements rather than the original measurements for state update. The differential measurements, i.e., the differences between the predicted distances and actual distance measurements, can be expressed as:
$\delta{d_{ij}^{\mathcal{V}}}$ $={\hat{d}_{ij}^{\mathcal{V}}}-{\tilde{d}_{ij}^{\mathcal{V}}}=\boldsymbol{H}_{ij}^{\mathcal{V}}\delta\boldsymbol{x}_{i}+\boldsymbol{G}_{ij}^{\mathcal{V}}\boldsymbol{v}_{ij}^{\mathcal{V}}$ (9)
@@
Here, $\boldsymbol{v}_{ij}^{\mathcal{V}}$ and $v_{im}^{\mathcal{A}}$ are the noise of the differential measurements, where $v_{im}^{\mathcal{A}}=-n_{d_{im}^{\mathcal{A}}}$, and the expressions of $\boldsymbol{v}_{ij}^{\mathcal{V}}$ and its corresponding covariance matrix $\boldsymbol{R}_{v_{ij}}$ are given by
$\boldsymbol{v}_{ij}^{\mathcal{V}}=\left[\left(\prescript{I}{}{\delta{\boldsymbol{p}}_{j}^{n}}\right)^{\mathsf{T}},\left(\boldsymbol{\phi}_{j}^{n}\right)^{\mathsf{T}},n_{d_{ij}^{\mathcal{V}}}\right]^{\mathsf{T}}$ (10)
$\boldsymbol{R}_{ij}^{\mathcal{V}}=\begin{bmatrix}&\boldsymbol{P}_{{p}_{j}},&\boldsymbol{P}_{{p}_{j}{\phi}_{j}},&\boldsymbol{0}_{3\times 1}\\ &\boldsymbol{P}_{{p}_{j}{\phi}_{j}}^{\mathsf{T}},&\boldsymbol{P}_{{\phi}_{j}},&\boldsymbol{0}_{3\times 1}\\ &\boldsymbol{0}_{1\times 3},&\boldsymbol{0}_{1\times 3},&\sigma_{r}^{2}\end{bmatrix}$ (11)
Here $\boldsymbol{P}_{{p}_{j}}$, $\boldsymbol{P}_{{\phi}_{j}}$, $\boldsymbol{P}_{{p}_{j}{\phi}_{j}}$ are the submatrices of the state covariance matrix of vehicle $j$, corresponding to the covariance matrix of its position, attitude, and their cross-covariance, respectively. Additionally, the expressions of $\boldsymbol{H}_{ij}^{\mathcal{V}}$, $\boldsymbol{G}_{ij}^{\mathcal{V}}$ and $\boldsymbol{H}_{ij}^{\mathcal{A}}$ are given by
$\boldsymbol{H}_{ij}^{\mathcal{V}}$ $=\left[\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}}\right)^{\mathsf{T}},\boldsymbol{0}_{3\times 3},\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}}\right)^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right],\boldsymbol{0}_{3\times 6}\right]$ (12)
$\boldsymbol{G}_{ij}^{\mathcal{V}}$ $=\left[-\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}}\right)^{\mathsf{T}},-\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}}\right)^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right],-1\right]$
@@
Assume that vehicle $i$ obtains the distances to $M$ anchors and $N$ other vehicles for state updates at time step $k$. Without loss of generality, let the indices of the $M$ anchors and $N$ other vehicles be $1$ to $M$ and $1$ to $N$, respectively. Let the measurement vector formed by the differential distances to the $M$ anchors be denoted as $\boldsymbol{z}_{i}^{\mathcal{A}}=\left[\delta d_{i1}^{\mathcal{A}},\cdots,\delta d_{iM}^{\mathcal{A}}\right]^{\mathsf{T}}\in\mathbb{R}^{M\times 1}$, with the corresponding covariance matrix being $\boldsymbol{R}_{i}^{\mathcal{A}}=\sigma_{r}^{2}\boldsymbol{I}_{M}$; Let the differential distances to the $N$ other vehicles and their corresponding covariance matrices be denoted as $\left\{\delta d_{ij}^{\mathcal{V}},\boldsymbol{R}_{ij}^{\mathcal{V}}\right\}$, $j=1,\ldots,N$. The predicted state and covariance matrix of vehicle $i$ before the update are given by $\left\{\hat{\boldsymbol{x}}_{i,k|k-1},\boldsymbol{P}_{i,k|k-1}\right\}$.
@@ III-B2 Independent Update
The differential distances between vehicle $i$ and anchors are independent of the vehicle’s state, and they can be directly used for state update via EKF. Since the distance measurements to the $M$ anchors are mutually independent, there is no difference between updating them sequentially as individual observations or combining them into a single observation vector for concurrent update. We choose the latter approach, and the corresponding update procedure can be written as:
@@ III-B3 Correlated Update
For distance measurements between vehicles, consider the measurement between vehicle $i$ and vehicle $j$ as an example, as shown in (9) and (10). The differential distance $\delta d_{ij}^{\mathcal{V}}$ is used for updating vehicle $i$’s state, and its noise includes the state estimation error of vehicle $j$. This leads to potential correlations between the differential distance measurement and vehicle $i$’s state. In a distributed framework, tracking the state correlation between vehicle $i$ and vehicle $j$ is extremely challenging or even infeasible. If the correlation is ignored and assumed to be independent, it could lead to the “data incest” problem [30], resulting in state estimation divergence. Therefore, our work focuses on this correlated update process, with the goal of achieving effective state update using relative distance measurements between vehicles under unknown correlations, ensuring accurate and consistent state estimation in 3D cooperative localization.
IV WCI-based Correlated Update
In this section, we present our methods for DCL in 3D scenarios. We begin by revisiting CI, including the version with weighting mechanism, WCI. Next, we propose a WCI-based approach for correlated updates in DCL, which accounts for the characteristics of CI fusion and the error propagation rule of INS, balancing the estimation of all state components with scale and observability disparities.
@@ IV-A Covariance Intersection Revisit
Typically, the cost function is chosen as the trace or determinant of the fused covariance matrix, i.e., $f(\boldsymbol{P})=\operatorname{tr}(\boldsymbol{P})$ or $\text{det}(\boldsymbol{P})$; we refer to them as the traditional criteria.
@@
The traditional trace-based criterion is derived from the principle of minimizing MSE of the estimate—equivalently, the trace of the actual covariance matrix (cf. (18)). In practice, since the actual covariance matrix is unavailable, the trace of the nominal covariance matrix is minimized instead.
@@
However, minimizing MSE alone is too coarse and lacks task specificity. To address this, Arambel et al. introduced relative weights to control the CI fusion process [19], termed WCI. However, they treated these weights merely as tuning knobs to improve the accuracy of a chosen component, without providing a theoretical rationale. We show in (19) that minimizing the weighted trace of the fused covariance matrix is mathematically equivalent to minimizing a WMSE over the entire state. By selecting the weighting matrix $\boldsymbol{W}$ to reflect task‐dependent priorities or target WMSE levels, WCI introduces task relevance and flexibility to the fusion criterion.
@@
In many applications, all state components are crucial and must be accounted for in the data fusion process. However, when state components differ significantly in units or magnitudes, the CI fusion process under traditional criteria tends to be dominated by a particular state component, neglecting the others. To demonstrate this, we present a simple two-dimensional example. Let $\boldsymbol{P}_{1}=\begin{bmatrix}1&0\\ 0&0.01\end{bmatrix}$ and $\boldsymbol{P}_{2}=\begin{bmatrix}1.09&0\\ 0&0.001\end{bmatrix}$. Although $\boldsymbol{P}_{2}$ reduces the second component’s variance by 90% dramatically, its 9% increase in the first component drives the trace-minimization fusion (CI-trace) to favor $\boldsymbol{P}_{1}$. As shown in Fig. 4, the fused covariance matrix thus remains almost identical to $\boldsymbol{P}_{1}$, offering negligible improvement for the small-scale component. By contrast, determinant-based criterion (CI-det) favors $\boldsymbol{P}_{2}$, as the the determinant corresponds to the uncertainty volume in which components with small scales can influence the estimation. However, both criteria lack the ability to flexibly adjust and balance the estimation performance of different components in the presence of significant scale disparity. By selecting an appropriate weighting matrix and applying WCI, it becomes possible to effectively balance the fusion process and fully leverage the information from both sources.
Figure 4: CI fusion under scale disparities (axes scaled differently for better demonstration). The covariance ellipse of CI-trace almost overlaps that of $\boldsymbol{P}_{1}$, indicating that the fusion is dominated by the larger-scale state component. As a result, the fused estimate closely follows the lower-uncertainty estimate along that dimension. While CI-det can account for smaller-scale state component, it still fails to achieve a proper balance between two components. In contrast, WCI effectively balances the fusion process and makes full use of the information from both sources.
IV-B Concurrent Update Strategy